# MagicDistillation: Weak-to-Strong Video Distillation for Large-Scale Portrait Few-Step Synthesis

### v1

@@ 3.2 Weak-to-Strong Distribution Matching
Figure 4: Vanilla DMD vs. W2SVD. Vanilla DMD encounters a significant challenge that $p_{\textrm{real}}$ has no overlap with the feasible region of the sample synthesized by the few-step generator $G_{\phi}$, which leads to an inaccurate estimation of the gradient difference $\nabla\log p_{\textrm{fake}}-\nabla\log p_{\textrm{real}}$. In contrast, W2SVD mitigates this issue by subtly shifting $p_{\textrm{real}}$ toward $p_{\textrm{fake}}$. This technique adjustment ensures a more substantial overlap between the two distributions, thereby enhancing the accuracy of the gradient estimation process.
@@
Table 1: Quantitative results comparison using HunyuanVideo (finetuned on portrait video synthesis task) on our customized VBench. All videos are 540$\times$960$\times$129 and are saved at 24fps. Each sample follow VBench (Huang et al., 2024a) to synthesize 5 videos to avoid errors.
@@
W2SVD w/o reg. loss (ours) 0.9763 0.9533 0.9954 0.7683 0.4899 0.5708 0.9993 0.8219
W2SVD w/ reg. loss (ours) 0.9735 0.9478 0.9955 0.0379 0.5732 0.7103 0.9939 0.7474
@@
W2SVD w/o reg. loss (ours) 0.9620 0.9656 0.9961 0.0523 0.5512 0.6909 0.9953 0.7448
Figure 5: Ablation studies of $\alpha_{\textrm{weak}}$ and $\mathcal{L}_{\textrm{reg}}$ using 4-step models on our customized VBench. W2SVD reduces to the vanilla DMD2 when $\alpha_{\textrm{weak}}$=0. From the average metrics presented in the lower right corner, it is evident that W2SVD achieves its optimal performance when $\alpha_{\textrm{weak}}$=0.25. Furthermore, the regularization loss (i.e., $\mathcal{L}_{\textrm{reg}}$) represents a trade-off between motion dynamics and visual quality.
@@
Figure 6: Visualization of different accelerated sampling methods (i.e., step distillation) on the HunyuanVideo model. In both 1-step and 4-step scenarios, W2SVD visibly outperforms other methods in terms of visual quality and facial motion dynamics.
Another advantage of incorporating LoRA is that W2S distribution matching can be achieved by adjusting the weight factor. As illustrated in Fig. 4, this method addresses the issue where fails to overlap $p_{\textrm{real}}$ with the feasible region of samples synthesized by the few-step generator $G_{\phi}$. To be specific, we observe that HunyuanVideo exhibits a large discrepancy between $p_{\textrm{real}}$ and $p_{\textrm{fake}}$ during DMD training for portrait video synthesis tasks, as illustrated on the left side of Fig. 4. This mismatch leads to an imprecise estimation of $\nabla\log p_{\textrm{fake}}-\nabla\log p_{\textrm{real}}$, ultimately synthesizing videos that lack controllability, as shown in Fig. 1. Intuitively, this issue typically emerges when the difference between $\Theta$ and $\theta$ surpasses the acceptable tolerance range. A practical solution to address this problem is to slightly adjust $p_{\textrm{real}}$ to better align it with $p_{\textrm{fake}}$. This adjustment can be effectively achieved by adjusting the LoRA weight factors $\alpha_{\textrm{weak}}$ (used for $\mathbf{v}_{\Theta}^{\textrm{real}}$) and $\alpha_{\textrm{strong}}$ (used for $\mathbf{v}_{\theta}^{\textrm{fake}}$):
@@
We further leverage Proposition 3.1 to demonstrate that W2S distribution matching not only ensures training stability but also achieves the same optimization objective as the standard distribution matching. Notably, as LoRA serves as an intermediate teacher, W2S distribution matching also enhances the performance of step distillation. Our empirical results (see Sec. 4) show that W2S distribution matching surpasses vanilla DMD in terms of motion dynamics, semantic faithfulness and visual quality.
@@ 3.3 Ground Truth Supervision
Another essential component of W2SVD is ground truth supervision, which is designed to enhance the stability of distribution matching while improving the visual quality of synthesized videos. The few-step generator is expected to match distributions not only between $p_{\textrm{real}}$ and $p_{\textrm{fake}}$ but also between $p_{\textrm{gt}}$ (i.e., $p(\mathbf{x}^{\textrm{gt}})$) and $p_{\textrm{fake}}$. This matching serves as a corrective measure against the challenges of unstable training process and the degradation of visual quality caused by the inherent biases present in $\mathbf{v}_{\Theta}^{\textrm{pre-train}}$. In our implementation. we minimize the following KL divergence $\mathcal{L}_{\textrm{reg}}$:
@@
where $\eta$ stands for a very small amount. As demonstrated in Sec. 4, we demonstrate that $\mathcal{L}_{\textrm{reg}}$ effectively facilitates a trade-off between motion dynamics and visual quality.
@@ 3.4 Detailed Implementation
Tailored Discriminator.
Adversarial training is a useful technique for enhancing the performance of step distillation. However, traditional discriminator architectures have predominantly relied on convolution-based designs. To better adapt to HunyuanVideo, we employ a DiT-based discriminator. We utilize the output of the final double block as the input to the discriminator and adopt the first layer of the single blocks, leveraging pre-training weights while ensuring no shared parameters, as the discriminator architecture.
1-Step Generator Training.
Using synthesized noise-data pairs for distillation (Liu et al., 2022) can significantly enhance the performance of the 1-step generator. However, synthesizing noise-data pairs using the 28-step HunyuanVideo model is computationally expensive and time-intensive. To mitigate this, we employ the 4-step generator obtained from W2SVD to synthesize noise-data pairs. Since the computational cost of 4-step sampling is relatively low, this process can be seamlessly integrated into the W2SVD training pipeline, eliminating the need for pre-synthesizing noise-data pairs. Initially, we set the distillation loss weight to 1 and the loss weight of $\mathcal{L}_{\textrm{DMD}}$ to 0.25. During the training of the 1-step generator, we progressively adjust the distillation (Liu et al., 2022) loss weight to 0.25 while increasing the loss weight of $\mathcal{L}_{\textrm{DMD}}$ to 1, which can ensure a balanced optimization process.
LoRA in HunyuanVideo.
We find that when utilizing bfloat16 precision within the deepspeed framework (Microsoft, 2022), the F-norm of the gradient during LoRA fine-tuning is significantly smaller compared to that observed during full-parameter fine-tuning. This discrepancy results in the vanishing gradient issue. Given the constraints imposed by using ZeRO3 in deepspeed, we address this problem by multiplying the loss associated with LoRA by an additional weight factor of 1e4, enabling stable and effective training under normal conditions.
@@ 4 Experiment
All experiments involving W2SVD, vanilla DMD/DMD2, and LCM for the 4-step generator were conducted using 8 NVIDIA H100 GPUs. For the 1-step generator, experiments with W2SVD and vanilla DMD/DMD2 were performed using 16 NVIDIA H100 GPUs. Our training data is using high quality widescreen videos from Youtube to match the resolution of 540$\times$960, and adding a portion of anime data to ensure W2SVD’s generalization ability. Our evaluation framework incorporates FID and FVD metrics assessed on the VFHQ (Xie et al., 2022), HDTF (Zhang et al., 2021), and Celeb-V (Zhu et al., 2022) datasets, alongside our customized VBench (Huang et al., 2024a, b), which includes seven dimensional metrics such as dynamic degree and image quality. Comprehensive details regarding the training hyperparameter configurations and the construction of the benchmarks are provided in Appendix B and A, respectively.
Since there is no open-source implementation of HunyuanVideo for portrait video synthesis, we first fine-tuned HunyuanVideo to derive the I2V-generator and TI2V-generator, both capable of synthesizing talking videos. It is important to note that our primary experiments were conducted with I2V-HunyuanVideo, while the experimental results for TI2V-HunyuanVideo are detailed in Appendix D.
4.1 Main Result
Our comparison experiments were conducted on our customized VBench, as well as on VFHQ, HDTF, and Celeb-V datasets. For VBench, we present the experimental results in Table 1. Here, the 28-step generator represents the fine-tuned HunyuanVideo model (i.e., the standard teacher model) specifically adapted for the portrait video synthesis task. Furthermore, the results of Euler (baseline) across all step scenarios are sampled using this fine-tuned HunyuanVideo model. For W2SVD, “w/o reg. loss” and “w/ reg. loss” denote not using ground truth supervision and using ground truth supervision, respectively. In the 4-step scenario, as shown in Table 1, it can be observed that W2SVD (w/o reg. loss) surpasses DMD2 across all metrics except imaging quality, with the average score exceeding DMD2 by a margin of 1.56%. Furthermore, W2SVD (w/o reg. loss) demonstrates a substantial improvement, achieving an increase of over 10% in the average score compared to both LCM and Euler. In particular, W2SVD (w/o reg. loss) even surpasses the 28-step generator (i.e., the teacher model) across all metrics except for imaging quality and aesthetic quality. We attribute this phenomenon to the ability of distribution matching to partially correct the intrinsic bias present in the fine-tuned HunyuanVideo model. Moreover, ground truth supervision facilitates a trade-off between visual quality and motion dynamics. Specifically, W2SVD (w/ reg. loss) reduces motion dynamics compared to W2SVD (w/o reg. loss) but achieves improvements in imaging quality and aesthetic quality. Note that although the motion dynamics of W2SVD (w/ reg. loss) are significantly lower than those of W2SVD (w/o reg. loss), the videos synthesized by W2SVD (w/ reg. loss) actually align more closely with the natural motion patterns of the human face when observed visually. In contrast, W2SVD (w/o reg. loss) still exhibits issues with instability in some synthesized videos on VBench. In the 1-step scenario, W2SVD (w/o reg. loss) surpasses DMD2 across all metrics except for i2v subject, subject consistency, and temporal flickering, achieving a 2.59% higher average score overall compared to DMD2.
Table 2: Quantitative results comparison on VFHQ. All videos are in the shape of 540$\times$960$\times$129 and are saved at 24fps.
@@ 4.1 Main Result
W2SVD w/o reg. loss (ours) 4 32.37 163.40
W2SVD w/ reg. loss (ours) 4 31.10 239.28
Table 3: Quantitative results comparison on HDTF. All videos are in the shape of 540$\times$960$\times$129 and are saved at 24fps.
@@
W2SVD w/o reg. loss (ours) 4 24.90 233.80
W2SVD w/ reg. loss (ours) 4 21.92 240.02
Table 4: Quantitative results comparison on Celeb-V. All videos are in the shape of 540$\times$960$\times$129 and are saved at 24fps.
@@
W2SVD w/o reg. loss (ours) 4 60.85 542.72
W2SVD w/ reg. loss (ours) 4 58.71 695.65
We report the FID and FVD evaluation results on the VFHQ, HDTF, and Celeb-V datasets in Tables 23, and 4, respectively. As demonstrated by these experimental results, W2SVD (w/o reg. loss) outperforms Euler, LCM, and DMD2 in the 4-step scenario across both FID and FVD metrics. Notably, across all datasets and metrics, W2SVD (w/o reg. loss) surpasses the 28-step teacher model, with the sole exception of the FID metric on the VFHQ dataset, where it slightly underperforms the 28-step teacher model. Additionally, incorporating ground truth supervision results in a further reduction in FID but leads to an increase in FVD, which we attribute to the fact that ground truth supervision negatively impacts motion dynamics. Consequently, by appropriately adjusting $\mathcal{L}_{\textrm{reg}}$ during the experiment, it is possible to achieve a desirable balance between motion dynamics and visual quality.

### v2

@@ 3.2 Weak-to-Strong Distribution Matching
Table 1: I2V quantitative results comparison using Magic141 (trained on both portrait and general video datasets) on our customized VBench. All videos are 540$\times$960$\times$129 and are saved at 24fps. Each sample follow VBench [10] to synthesize 5 videos to avoid errors.
@@
MagicDistillation w/o reg. loss 0.9763 0.9533 0.9954 0.7683 0.4899 0.5708 0.9993 0.8219
MagicDistillation w/ reg. loss 0.9735 0.9478 0.9955 0.0379 0.5732 0.7103 0.9939 0.7474
@@
MagicDistillation w/o reg. loss 0.9620 0.9656 0.9961 0.0523 0.5512 0.6909 0.9953 0.7448
Figure 4: Ablation studies of $\alpha_{\textrm{weak}}$ and $\mathcal{L}_{\textrm{reg}}$ using 4-step models on our customized VBench. MagicDistillation reduces to the vanilla DMD2 when $\alpha_{\textrm{weak}}$=0. From the average metrics presented in the lower right corner, it is evident that MagicDistillation achieves its optimal performance when $\alpha_{\textrm{weak}}$=0.25. Furthermore, when the visual quality of the ground truth video is high but the motion dynamic is insufficient, the regularization loss (i.e., $\mathcal{L}_{\textrm{reg}}$) represents a trade-off between motion dynamics and visual quality.
Figure 5: Vanilla DMD vs. MagicDistillation. Vanilla DMD encounters a significant challenge that $p_{\textrm{real}}$ has no overlap with the feasible region of the sample synthesized by the few-step generator $G_{\phi}$, which leads to an inaccurate estimation of the gradient difference $\nabla\log p_{\textrm{fake}}-\nabla\log p_{\textrm{real}}$. In contrast, MagicDistillation mitigates this issue by subtly shifting $p_{\textrm{real}}$ toward $p_{\textrm{fake}}$. This technique adjustment ensures a more substantial overlap between the two distributions, thereby enhancing the accuracy of the gradient estimation process.
@@
Table 2: TI2V quantitative results comparison between MagicDistillation and other populer large-scale VDMs on our customized VBench. In particular, the VDM referenced for 28-step sampling using the Euler (baseline) in Table 1 corresponds to Magic141 as presented in this table.
50-step
  WanX-I2V (14B) 0.9800 0.9649 0.9910 0.5625 0.5980 0.7031 0.9858 0.8264
HunyuanVideo-I2V (13B) 0.9834 0.9654 0.9959 0.1639 0.5992 0.6990 0.9940 0.7715
Magic141 (13B) 0.9677 0.9607 0.9936 0.7704 0.5892 0.6906 0.9895 0.8516
4-step
  MagicDistillation 0.9659 0.9374 0.9881 0.9935 0.5782 0.7122 0.9727 0.8782
Figure 6: Visualization of various accelerated sampling methods under the 4-step scenario. Additional visualizations for the 1-step scenario and other details are provided in Appendix I.
Another advantage of incorporating LoRA is that W2S distribution matching can be achieved by adjusting the weight factor. As illustrated in Fig. 5, this method addresses the issue where fails to overlap $p_{\textrm{real}}$ with the feasible region of samples synthesized by the few-step generator $G_{\phi}$. To be specific, we observe that Magic141 exhibits a large discrepancy between $p_{\textrm{real}}$ and $p_{\textrm{fake}}$ during DMD training for portrait video synthesis tasks, as illustrated on the left side of Fig. 5. This mismatch leads to an imprecise estimation of $\nabla\log p_{\textrm{fake}}-\nabla\log p_{\textrm{real}}$, ultimately synthesizing videos that lack controllability, as shown in Fig. 2 (Left). Intuitively, this issue typically emerges when the difference between $\Theta$ and $\theta$ surpasses the acceptable tolerance range. A practical solution to address this problem is to slightly adjust $p_{\textrm{real}}$ to better align it with $p_{\textrm{fake}}$. This adjustment can be effectively achieved by adjusting the LoRA weight factors $\alpha_{\textrm{weak}}$ (used for $\mathbf{v}_{\Theta}^{\textrm{real}}$) and $\alpha_{\textrm{strong}}$ (used for $\mathbf{v}_{\theta}^{\textrm{fake}}$):
@@
We further leverage Proposition 3.1 to demonstrate that W2S distribution matching not only ensures training stability but also achieves the same optimization objective as the standard distribution matching. Notably, as LoRA serves as an intermediate teacher, W2S distribution matching also enhances the performance of step distillation. Our empirical results (see Sec. 4) show that W2S distribution matching surpasses vanilla DMD in terms of motion dynamics, semantic faithfulness and visual quality. By utilizing higher-quality portrait and general video datasets for step distillation, MagicDistillation, with a mere 4-step sampling process (i.e., NFE$=$4), is capable of outperforming WanX-I2V, HunyuanVideo-I2V, and Magic141 in portrait video synthesis, even under their 50-step sampling scenarios (i.e., NFE$=$100).
@@ 3.3 Ground Truth Supervision
Another essential component of MagicDistillation is ground truth supervision, which is designed to enhance the stability of distribution matching while improving the visual quality of synthesized videos. The few-step generator is expected to match distributions not only between $p_{\textrm{real}}$ and $p_{\textrm{fake}}$ but also between $p_{\textrm{gt}}$ (i.e., $p(\mathbf{x}^{\textrm{gt}})$) and $p_{\textrm{fake}}$. This matching serves as a corrective measure against the challenges of unstable training process and the degradation of visual quality caused by the inherent biases present in $\mathbf{v}_{\Theta}^{\textrm{pre-train}}$. In our implementation. we minimize the following KL divergence $\mathcal{L}_{\textrm{reg}}$:
@@
where $\eta$ stands for a very small amount. As demonstrated in Sec. 4, we demonstrate that $\mathcal{L}_{\textrm{reg}}$ is effective in pulling together the distribution of the synthesized video obtained from few-step generator and the distribution of the ground truth video.
@@ 3.4 Detailed Implementation
Tailored Discriminator. Adversarial training is a useful technique for enhancing the performance of step distillation. However, traditional discriminator architectures have predominantly relied on convolution-based designs. It is noteworthy that the backbone architecture of Magic141 consists of multi-model (MM) DiT blocks, which serves as the foundation for the double blocks in Magic141, alongside vanilla DiT blocks, which forms the single blocks in the corresponding counterpart. To better align with the structure of Magic141, we incorporate a DiT-based discriminator. Specifically, we use the output from the final double block as the input for the discriminator and employ the initial layer of the single blocks, capitalizing on pre-trained weights while ensuring that no parameters are shared, as the foundation for the discriminator’s architecture.
1-Step Generator Training. Using synthesized noise-data pairs for distillation [17] can significantly enhance the performance of the 1-step generator. However, synthesizing noise-data pairs using Magic141 with 50 sampling steps is computationally expensive and time-intensive. To mitigate this, we employ the 4-step generator obtained from MagicDistillation to synthesize noise-data pairs. Since the computational cost of 4-step sampling is relatively low, this process can be seamlessly integrated into the MagicDistillation training pipeline, eliminating the need for pre-synthesizing noise-data pairs. Initially, we set the distillation loss weight to 1 and the loss weight of $\mathcal{L}_{\textrm{DMD}}$ to 0.25. During the training of the 1-step generator, we progressively adjust the distillation [17] loss weight to 0.25 while increasing the loss weight of $\mathcal{L}_{\textrm{DMD}}$ to 1, which can ensure a balanced optimization process.
LoRA in Fake DiT and Real DiT. We find that when utilizing bfloat16 precision within the deepspeed framework [19], the F-norm of the gradient during LoRA fine-tuning is significantly smaller compared to that observed during full-parameter fine-tuning. This discrepancy results in the vanishing gradient issue. Given the constraints imposed by using ZeRO3 in deepspeed, we address this problem by multiplying the loss associated with LoRA by an additional weight factor of 1e4, enabling stable and effective training under normal conditions.
@@ 4 Experiment
All experiments involving MagicDistillation, vanilla DMD/DMD2, and LCM for the 4-step generator, with the exception of those detailed in Table 2, Table 3 and Table 7, were carried out using 8 NVIDIA H100 GPUs. In addition to this, MagicDistillation in Table 2, Table 3 and Table 7 uses 40 NVIDIA H100 GPUs. For the 1-step generator, experiments with MagicDistillation and vanilla DMD/DMD2 were performed using 16 NVIDIA H100 GPUs. For all datasets except those used for MagicDistillation in Table 2, our training data comprises high-quality widescreen videos sourced from YouTube, carefully curated to match the resolution of 540$\times$960. Additionally, we incorporate a selection of anime data to further enhance the generalization capability of MagicDistillation. For the experimental results showcased in Table 2, we incorporate a selected portion of general data. This data is carefully curated from Koala-36M [35] and Intern4k [28], with stringent conditions to ensure both superior aesthetic quality and dynamic motion attributes.
@@
Our evaluation framework incorporates FID and FVD metrics assessed on the VFHQ [37], HDTF [49], and Celeb-V [51] datasets, the general I2V-VBench [11], alongside our customized VBench [10, 11], which includes seven dimensional metrics such as dynamic degree and image quality. Comprehensive details regarding the training hyperparameter configurations and the construction of the benchmarks are provided in Appendix B and A, respectively.
We employ Magic141 [40] as the teacher model to perform step distillation. Magic141, a 13B open-source large-scale VDM, is developed through the co-training of portrait video and general video datasets. This model serves as both our I2V-teacher and TI2V-teacher. For I2V synthesis, we simplify the prompt by setting it to null during model training and performance evaluation. It is worth noting that the teacher model referenced in Tables 166 and 6, along with the few-step generators such as DMD/DMD2, LCM, and MagicDistillation, are exclusively I2V-generators. Besides, the models referenced in the remaining tables are all TI2V-generators. Further experimental results pertaining to the TI2V-generator are comprehensively detailed in Appendix E.
Table 3: TI2V quantitative results comparison between MagicDistillation and other popular large-scale VDMs on general I2V-VBench. Experimental results for the other comparison methods were directly extracted from the official VBench website. Note that the results for WanX-I2V and HunyuanVideo-I2V are not presented here, as the official website does not provide data for these VDMs.
Method i2v subject (↑) subject consistency (↑) motion smoothness (↑) dynamic degree (↑) aesthetic quality (↑) imaging quality (↑) average (↑)
100-NFE
  CogVideoX-I2V-SAT [39] 0.9767 0.9547 0.9835 0.3651 0.5976 0.6764 0.7590
I2Vgen-XL [47] 0.9752 0.9636 0.9831 0.2496 0.6533 0.6985 0.7538
SEINE-512x320 [3] 0.9657 0.9420 0.9668 0.3431 0.5842 0.7097 0.7519
VideoCrafter-I2V [2] 0.9117 0.9786 0.9800 0.2260 0.6078 0.7168 0.7368
SVD-XT-1.0 [1] 0.9752 0.9552 0.9809 0.5236 0.6015 0.6980 0.7890
Step-Video-TI2V [18] 0.9550 0.9602 0.9924 0.4878 0.6299 0.7044 0.7882
4-NFE
  MagicDistillation 0.9512 0.9006 0.9766 0.9840 0.5955 0.6949 0.8505
Table 4: I2V comparison on VFHQ. Videos are in the shape of 540$\times$960$\times$129.
@@
MagicDistillation w/o reg. loss 4 32.37 163.40
MagicDistillation w/ reg. loss 4 31.10 239.28
Table 5: I2V comparison on HDTF. Videos are in the shape of 540$\times$960$\times$129.
@@
MagicDistillation w/o reg. loss 4 24.90 233.80
MagicDistillation w/ reg. loss 4 21.92 240.02
Table 6: I2V comparison on Celeb-V. Videos are in the shape of 540$\times$960$\times$129.
@@
MagicDistillation w/o reg. loss 4 60.85 542.72
MagicDistillation w/ reg. loss 4 58.71 695.65
4.1 Main Result
Our comparison experiments were conducted on our customized VBench, as well as on VFHQ, HDTF, and Celeb-V datasets. For VBench, we present the experimental results in Table 1. Here, the 28-step generator represents the pre-trained Magic141 (i.e., the standard teacher model) specifically adapted for the portrait video synthesis task. Furthermore, the results of Euler (baseline) across all step scenarios are sampled using the pre-trained Magic141. For MagicDistillation, “w/o reg. loss” and “w/ reg. loss” denote not using ground truth supervision and using ground truth supervision, respectively. In the 4-step scenario, as shown in Table 1, it can be observed that MagicDistillation (w/o reg. loss) surpasses DMD2 across all metrics except imaging quality, with the average score exceeding DMD2 by a margin of 1.56%. Furthermore, MagicDistillation (w/o reg. loss) demonstrates a substantial improvement, achieving an increase of over 10% in the average score compared to both LCM and Euler. In particular, MagicDistillation (w/o reg. loss) even surpasses the 28-step generator (i.e., the teacher model) across all metrics except for imaging quality and aesthetic quality. We attribute this phenomenon to the ability of distribution matching to partially correct the intrinsic bias present in Magic141. Moreover, ground truth supervision can further enhance the visual quality of the synthesized video. Specifically, in Table 1, MagicDistillation (w/ reg. loss) achieves improvements in imaging quality and aesthetic quality compared to MagicDistillation (w/o reg. loss). Even while the motion dynamics of MagicDistillation (w/ reg. loss) are noticeably lower compared to MagicDistillation (w/o reg. loss), this outcome is primarily attributed to the training data consisting solely of talking videos, which inherently exhibit low motion dynamics. However, when the training data is supplemented with a mix of general videos featuring significantly more pronounced object motion, the motion degree of MagicDistillation, as presented in Table 2, reaches an impressive 0.9935. This value is significantly higher than other compared VDMs. In the 1-step scenario, MagicDistillation surpasses DMD2 across all metrics except for i2v subject, subject consistency, and temporal flickering, achieving a 2.59% higher average score overall compared to DMD2.
We further compare MagicDistillation, trained with higher-quality general videos, against WanX-I2V, HunyuanVideo-I2V, and Magic141, with the results detailed in Table 2 and Table 3. Remarkably, MagicDistillation achieves an outstanding score of 0.8782 on our customized VBench, outperforming all competing methods. Besides, MagicDistillation demonstrates exceptional performance on the general I2V-VBench, as substantiated by the results showcased in Table 3. Its motion degree achieves an impressive 0.9840, significantly surpassing both CogVideoX-I2V-SAT [38] and Step-Video-TI2V [18]. Furthermore, MagicDistillation attains a remarkable final average score of 0.8505, comfortably outperforming the other comparison VDMs, all of which fail to exceed a score of 0.8. The visualized results on our project page further underscore this achievement. Based on these findings, we conclude that even the most renowned open-source VDMs still exhibit notable limitations in portrait video synthesis.
We report the FID and FVD evaluation results on the VFHQ, HDTF, and Celeb-V datasets in Tables 66, and 6, respectively. As demonstrated by these experimental results, MagicDistillation (w/o reg. loss) outperforms Euler, LCM, and DMD2 in the 4-step scenario across both FID and FVD metrics. Notably, across all datasets and metrics, MagicDistillation (w/o reg. loss) surpasses the 28-step teacher model, with the sole exception of the FID metric on the VFHQ dataset, where it slightly underperforms the 28-step teacher model. Additionally, incorporating ground-truth supervision results in a further reduction in FID but leads to an increase in FVD, which we attribute to the negative impact of ground-truth supervision on motion dynamics when the training data’s motion dynamics are insufficient.